Medication Calculation Worksheeet #2 1. Ordered: Demerol 35 mg IM Available: Demerol 50 mg/ml Give:_________ 2. Ordered: Nembutal 100 mg capsule Available: Nembutal 0.1 gm/capsule Give:_________ 3. Ordered: Kannamycin 200 mg Available: Kannamycin 0.5 gm/2 ml Give:_________ 4. Ordered: Atropine 0.3 mg IV Stat Available: Atropine gr 1/150 per ml Give:_________ 5. Ordered: Ativan 0.5 mg IM PRN anxiety Available: Ativan 2 mg/ml Give:_________ 6. Ordered: Aspirin gr X PO Available: Aspirin 0.3 gm Give:_________ By what route? ___________ 7. Ordered: Premarin 1.25 mg PO QD Available: Premarin 0.65 mg/tablet Give:_________ 8. Ordered: Lithium 12 mcg PO BID Available: Lithium 8 mcg/ 5 ml Give:_________ 9. Ordered: Heparin 15,000 units SC Q 8 H Available: Heparin 20,000 units/ml Give:_________ in what syringe?________ 10. Ordered: Digoxin 0.125 mg PO QD Available: Digoxin 0.25 mg/tablet SHOW WORK
Answer #1

ANSWER.

1) Calculation of volume of drug required.

        Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

Here

     Desired dose(D) = 35 mg.

     Dose in hand(H) = 50 mg.

            Quantity(Q) = 1 ml.

As per above formula

     Volume(ml) of drug required = D/H *Q

                                                   = 35/50 * 1

                                                   = 0.7 ml.

2) Calculation of number of tablets required.

         Ordered dose = 100 mg.

         Available dose = 100 mg (0.1g)

    Number of tablet required = 100/100

                                              = 1 capsule.

3) Calculation of volume(ml) of drug required.

        Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

   Here

      Desired dose(D) = 200 mg.

      Dose in hand(H) = 500 mg.

           Quantity(Q) = 2 ml.

As per above formula

     Volume(ml) of drug required = D/H * Q

                                                   = 200/500 * 2

                                                   = 0.4 * 2

                                                   = 0.8 ml.

4) Calculation of volume of drug required.

           Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

    Here

       Desired dose(D) = 0.3 mg.

       Dose in hand(H) = gr 1/150.

Convert gr1/150 to mg for calculation

         1 gr = 60 mg

   Dose in hand (H) = 60/150

                               = 0.4 mg

        Quantity(Q) = 1 ml.

As per above formula

       Volume(ml) of drug required = D/H * Q

                                                     = 0.3/0.4 * 1

                                                     = 0.75 * 1

                                                     = 0.8 ml.

5) Calculation of volume of drug required.

          Volume = Desired dose(D)/Dose in hand(H) * Quantity(Q).

    Here

         Desired dose(D) = 0.5 mg.

         Dose in hand(H) = 2 mg.

            Quantity(Q) = 1 ml

As per above formula

     Volume(ml) of drug required = D/H * Q

                                                   = 0.5/2 * 1

                                                   = 0.25 ml.

6) Calculation of number of tablet required.

           Ordered dose = 600 mg (gr X = 600 mg).

           Available dose = 0.3 gm.

   Convert gm to mg for calculation

           1 gm = 1000 mg.

   So

     Available dose = 03 * 1000

                             = 300 mg.

      Number of tablets required = Ordered dose/Available dose

                                                  = 600/300

                                                  = 2 tablets by oral route.

7) Calculation of number of tablets required.

          Ordered dose = 1.25 mg.

           Available dose = 0.65 mg/tablet.

    Number of tablet required = Ordered dose/Available dose

                                              = 1.25/0.65

                                              = 1.9

                                              = 2 tablets.

8) Calculation of volume of drug required.

            Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

       Here

            Desired dose(D) = 12 mcg.

            Dose in hand(H) = 8 mcg.

                 Quantity(Q) = 5ml.

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 12/8 * 5

                                                    = 1.5 * 5

                                                   = 7.5 ml.

9) Calculation of volume of drug required.

          Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

      Here

          Desired dose(D) = 15000 unit.

          Dose in hand(H) = 20000 unit.

              Quantity(Q) = 1 ml.

As per above formula

          Volume(ml) of drug required = D/H * Q

                                                        = 15000/20000 * 1

                                                        = 0.75 ml using TB syringe.

10) Calculation of number of tablets required.

              Ordered dose = 0.125 mg.

              Available dose = 0.25 mg.

     Number of tablets required = Ordered dose/Available dose

                                                 = 0.125/0.25

                                                 = 0.5 (half of a tablet).